i need to simplify the expression (x^8y^-8/x^2y^-2)^1/2 i tried to solve it on my own but fractional exponents really throw me off...=/
Answer by richwmiller(17219)

(x^8y^-8/x^2y^-2)^1/2 means sqrt(x^8y^-8/x^2y^-2)=x^3/y^5
y^-8 means 1/y^8
y^-2=1/y^2
I did this same problem for you a couple days ago.